# Seg-Zero: Reasoning-Chain Guided Segmentation via Cognitive Reinforcement

### v1

@@ 4.4 Comparison with Other Methods
Reasoning Segmentation. We compare the zero-shot performance on ReasonSeg [17], results are shown in Table 6. We can find our Seg-Zero achieves the SOTA zero-shot performance acrosss various methods.
Referring Expression Segmentation. The results on referring expression segmentation is shown on Table 7. Moreover, we find that the ground-truth annotations in RefCOCO(+/g) are not precise enough, which suggests that our Seg-Zero model should, in principle, achieve better performance than values in the table. Supplementary materials provide detailed analysis.
@@
4.5 Qualitative Results
We provide several examples in Figure 6 and Figure 9. We can easily obeserve that the reasoning process is helpful in analyzing user instructions, especially when there are multiple objects within the same class categories. For instance, Seg-Zero demonstrates its ability to discern that the a ‘recreational vehicle’ is more appropriate than a ‘truck’ in the context of a ‘road trip’, and correctly indentifies that a ‘conductor’ is ‘positioned at the front of the stage’.
5 Conclusion
In this paper, we propose Seg-Zero, an novel framework that integrates CoT reasoning process into segmentation tasks. We design a sophisticated reward mechanism, incorporating both format and accuracy constraints, to guide the optimization directions. By training exclusively with RL, Seg-Zero emerges reasoning capabilities without relying on any supervised reasoning data. We present a detailed comparison between SFT and RL, as well as the introduction of reason chain. Additionally, we offer insightful perspectives on the design of RL and the reward functions.

### v2

@@ 4.4 Comparison with Other Methods
Reasoning Segmentation. We compare the zero-shot performance on ReasonSeg [17], results are shown in Table 9. We can find our Seg-Zero achieves the SOTA zero-shot performance across various methods.
Referring Expression Segmentation. The results on referring expression segmentation are shown on Table 10. Moreover, we find that the ground-truth annotations in RefCOCO(+/g) are not precise enough, which suggests that our Seg-Zero model should, in principle, achieve better performance than values in the table. Supplementary materials provide detailed analysis.
4.5 Qualitative Results
We provide several examples in Figure 7. We can easily observe that the reasoning process is helpful in analyzing user instructions, especially when there are multiple objects within the same class categories. For instance, Seg-Zero demonstrates its ability to discern that a ‘recreational vehicle’ is more appropriate than a ‘truck’ in the context of a ‘road trip’, and correctly identifies that a ‘conductor’ is ‘positioned at the front of the stage’.
5 Conclusion
In this paper, we propose Seg-Zero, a novel framework that integrates the CoT reasoning process into segmentation tasks. We design a sophisticated reward mechanism, incorporating both format and accuracy constraints, to guide the optimization directions. By training exclusively with RL, Seg-Zero emerges reasoning capabilities without relying on any supervised reasoning data. We present a detailed comparison between SFT and RL, as well as the introduction of reason chain. Additionally, we offer insightful perspectives on the design of RL and the reward functions.
Appendix A Rewards Changes During Training
Appendix B More Visualization Examples
We provide more examples in Figure 9.
Appendix C Additional Analysis of RefCOCO(+/g)
Through an analysis of the RefCOCO(+/g) datasets, we observe that the mask annotations in these datasets lack precision. Specifically, these annotations either inaccurately represent object sizes or exhibit imprecise edge handling, as illustrated in Figure 10. We randomly select 100 samples, and find that 5% samples has ‘inaccurately represent object sizes ’, and almost 70% samples is not precious in ‘edge hanlding’. Although our model demonstrates high precious, these annotation errors result in a lower IoU score.
Since previous methods [17, 7] fine-tunes the segmentation decoder on these datasets, they can adapt to these imprecise annotations. In contrast, our segmentation model remains frozen and maintains higher precision. Moreover, since the gIoU computes the aggregate of this bias, it fails to accurately reflect the true performance of our model. Therefore, the actual precious of our Seg-Zero should be better than the performance presented in Table 10. To further evaluate the localization accuracy of our Seg-Zero model, we introduce two additional metrics: bbox accuracy and point accuracy. The bbox accuracy metric computes whether the IoU between the predicted bounding box and the maximum bounding box that fully encloses the ground-truth mask exceeds 0.5. Meanwhile, the point accuracy metric assesses whether the predicted points lie within the ground-truth mask. These metrics provide insights into the model’s ability to precisely localize the target object. As shown in Table 11, our Seg-Zero model accurately localizes the target object with a success rate of approximately 90%.
Method ReasonSeg
val test
bbox point bbox point
Seg-Zero-3B 85.4 89.3 85.5 88.8
Seg-Zero-7B 85.4 89.0 86.6 90.0
Table 11: Bbox IoU and Point Accuracy.
@@ Appendix C Additional Analysis of RefCOCO(+/g)
Figure 10: Visual analysis of RefCOCO(+/g). The grouth-truth label on RefCOCO(+/g) is not precious, whereas our results are precious.
Appendix D Limitations and Future Works
Currently, our Seg-Zero only focuses on single object reasoning segmentation. We are considering employ our method to multi-object reasoning segmentation, instance-level reasoning segmentation.